# MA-RAG: Multi-Agent Retrieval-Augmented Generation via Collaborative Chain-of-Thought Reasoning

### v1

@@ 5 Conclusion
We present MA-RAG, a novel multi-agent framework for Retrieval-Augmented Generation that recasts the RAG pipeline as a coordinated reasoning process among specialized agents. Each agent, responsible for planning, retrieval, extraction, or generation, uses chain-of-thought prompting to enable structured and context-sensitive query resolution. Unlike prior work that tunes individual components or requires supervised training, MA-RAG is entirely training-free and generalizes well across domains and question types.
Empirical results on five open-domain and multi-hop QA benchmarks show that MA-RAG outperforms strong LLMs and state-of-the-art RAG baselines, achieving new best results on several datasets. Ablation studies confirm the importance of the planner and extractor: the former decomposes complex queries, while the latter improves retrieval precision. Strategic allocation of model capacity across agents yields further gains in both performance and efficiency.
Together, these findings highlight the potential of modular, agent-based reasoning as a scalable and adaptable approach to improving retrieval-augmented generation.
Appendix A Details of Experimental Dataset
A.1 Main Experiments
@@ Appendix B Experiment Results
Appendix C Implementation Details
@@
C.1 Workflows
We implement MA-RAG using LangChain and LangGraph [^2] https://www.langchain.com/, where agents communicate through structured JSON messages. Each agent is represented as a node in the graph, and edges determine the next agent to execute based on the current state. The overall graph structure representing the architecture of MA-RAG is shown in Figure LABEL:fig:langchain. For modularity and clarity, we define separate subgraphs for core components, including the planner, the plan executor, and the RAG pipeline.
C.2 Agent Communication
@@ Appendix D Boarder Impacts
Appendix E Prompt Formats
E.1 Planner Agent
System: You are tasked with assisting users in generating structured plans
for answering questions. Your goal is to deconstruct a query into
manageable, simpler components. For each question, perform these tasks:
*Analysis: Identify the core components of the question, emphasizing the
key elements and context needed for a comprehensive understanding.
Determine whether the question is straightforward or requires
multiple steps to provide an accurate answer.
*Plan Creation:
- Break down the question into smaller, simpler questions by reasoning
that lead to the final answer. Ensure those steps are non overlap.
- Ensure each step is clear and logically sequenced.
- Each step is a question to search, or to aggregate output from  previous
steps. Do not verify previous step.
# Notes:
- Put your output in a list of string, each string describe a sub-task
User: {Question}
E.2 Step Definer Agent
System: Given a plan, the current step, and the results from finished
steps, decide the task for this step.Output the type of task and the
query. The query need to be in detail, include all of information from
previous step’s results in the query if it maked, especially for aggregate
task. Be concise.
User:
Plan: {plan}
Current step: {cur_step}
Results of finished steps:
{memory}
E.3 Extractor Agent
System: Summarize and extract all relevant information from the provided
passages based on the given question. Remove all irrelevant information.
Think step-by-step.
**Identify Key Elements**: Read the question carefully to determine
what specific information is being requested.
**Analyze Passages**: Review the passages thoroughly to find any
segments that contain information relevant to the question.
**Extract Relevant Information**: Highlight or note down sentences,
phrases, or words from the passages that relate to the question.
**Remove Irrelevant Details**: Ensure that all extracted information
is relevant to the question, eliminating any unnecessary or unrelated
content.
# Output Format
- Output a list of notes. Each note contains related information from
the passage, and each note is clear, standalone.
# Notes
- Avoiding any irrelevant details.
- If a piece of information is mentioned in multiple places,
include it only once.
- If there are no related information, output:
No related information from this document.
User:
Passage: {passage}
Query: {question}?
E.4 Question-Answering Agent
System: You are an assistant for question-answering tasks. Use the
following process to deliver concise and precise answers based on
the retrieved context.
1. Analyze Carefully: Begin by thoroughly analyzing both the question
and the provided context.
2. Identify Core Details**: Focus on identifying the essential names,
terms, or details that directly answer the question. Disregard any
irrelevant information.
3. Provide a Concise Answer:
- Remove redundant words and extraneous details.
- Present the answer by listing only the necessary names, terms, or very
brief facts that are crucial for answering the question.
4. Clarity and Accuracy: Ensure that your answer is clear and maintains
the original meaning of the information provided.
5. Consensus: If the contexts are not consensus, pick one which is the
most logical, consensus, or confident.
User:
Retrieved information:
{context}
Question:
{question}
Appendix F Case Study
F.1 HotpotQA
@@
Q: What role did Thomas Doherty play in sitcom based on the Israeli series "North Star"? Correct Answer: Sean
MA-RAG 70B Plan: ["Identify the sitcom based on the Israeli series North Star.", "Determine the character role played by Thomas Doherty in that sitcom."]
@@
Extracted Information: doc-19587422: [No related information from this document.], doc-19587424: [No related information from this document.], doc-14539692: [No related information from this document.], doc-19587423: [The sitcom based on the Israeli series ’North Star’ is titled "The Lodge." This is evidenced by the passage stating, "The British Disney Channel produces an adapted under the title ’The Lodge’ (working title: ’North Star’)".], doc-17716638: [No related information from this document.]
@@
Llama3-70B Answer: Thomas Doherty played the role of Jasper in the sitcom "High Fidelity", which is based on the Israeli series "North Star."
ChatQA 70B Query: What role did Thomas Doherty play in sitcom based on the Israeli series "North Star"?
@@
Doc-20322850: Thomas Doherty (actor) Thomas Doherty (born 21 April 1995) is a Scottish actor, best known for his roles as Harry Hook in Disney Channel’s "Descendants 2" and as Sean in the British musical comedy-drama, "The Lodge". Doherty was born and raised in Edinburgh, Scotland. He has an older brother and younger sister, and his parents were both in the banking business. Doherty began acting at the age of 5 and also played football. Doherty joined Lothian Youth Arts and Musicals Company and played a variety of roles, including his debut as Disney’s Aladdin. Doherty attended Royal High School in Edinburgh.
@@
Doc-12942841: them. The show was canceled after one season owing to poor ratings. She also starred in the popular British sitcom ""Bo! in the USA"", a brain child of Leigh Francis. In the show, she plays herself being randomly harassed by Avid Merrion (Francis), who claims they are lovers. The show aired in October 2006 on the British Channel 4, and she appeared in several episodes. During 2007, Doherty appeared in two television projects. She first appeared in ""Kiss Me Deadly: A Jacob Keane Assignment"" for the Here TV network and followed up with a starring role in the holiday film
Doc-1423761: to the stage, making a disastrous start in John Phillips’ ""Man on the Moon"" (1975). In 1977, he returned to his birthplace, Halifax, Nova Scotia, and started playing legitimate roles, including Shakespeare, at the Neptune Theatre under the tutelage of John Neville. This led to television work, beginning with a variety program, ""Denny’s Sho*"", which ran for one season in 1978. He went on to host and voice parts in the children’s program, ""Theodore Tugboat"", and to act in various series, including twenty-two episodes of the drama ""Pit Pony"". Doherty also performed with the New Mamas and the Papas (see
@@
ReAct (70B) Query: What role did Thomas Doherty play in sitcom based on the Israeli series "North Star"?
AI Messages: Retrieve(query="Thomas Doherty sitcom North Star role")
@@
Doc-20322850: Thomas Doherty (actor) Thomas Doherty (born 21 April 1995) is a Scottish actor, best known for his roles as Harry Hook in Disney Channel’s "Descendants 2" and as Sean in the British musical comedy-drama, "The Lodge". Doherty was born and raised in Edinburgh, Scotland. He has an older brother and younger sister, and his parents were both in the banking business. Doherty began acting at the age of 5 and also played football. Doherty joined Lothian Youth Arts and Musicals Company and played a variety of roles, including his debut as Disney’s Aladdin. Doherty attended Royal High School in Edinburgh.
Doc-1423761: to the stage, making a disastrous start in John Phillips"́"Man on the Moon"" (1975). In 1977, he returned to his birthplace, Halifax, Nova Scotia, and started playing legitimate roles, including Shakespeare, at the Neptune Theatre under the tutelage of John Neville. This led to television work, beginning with a variety program, ""Dennyś Sho*"", which ran for one season in 1978. He went on to host and voice parts in the childrenś program, ""Theodore Tugboat"", and to act in various series, including twenty-two episodes of the drama ""Pit Pony"". Doherty also performed with the New Mamas and the Papas (see
Doc-18484135: After high school, Doherty attended The MGA Academy of Performing Arts in Edinburgh, Scotland. He graduated after studying music theatre at MGA Academy in July 2015 and immediately signed a contract with Olivia Bell Management in London. Doherty was trained in acting, singing, and various types of dance including contemporary, hip hop, jazz, tap and ballet. After graduating from The MGA Academy of Performing Arts in the summer of 2015, Doherty began waiting tables in Edinburgh at Tigerlilly restaurant. On his days off, Doherty split his time between auditioning for ""The Lodge"" and working with the Edinburgh Fringe. Doherty was
@@
Doc-2573069: of Harbour Master, as well as the voice-overs of the characters, in ""Theodore Tugboat"", a CBC Television children’s show chronicling the ""lives"" of vessels in a busy harbour loosely based upon Halifax Harbour. In 1999, he also played Charley McGinnis in 22 episodes of the CBC Television series ""Pit Pony"". In 2004, Doherty appeared on Sharon, Lois Bram’s 25th Anniversary Concert special, ""25 Years of Skinnamarink"", that aired on CBC on January 1, 2004. He sang two songs with the trio: "California Dreamin" and "Who Put the Bomp?" One of his last appearances was in the Canadian TV series
Answer: Thomas Doherty played the role of Brandon in the sitcom based on the Israeli series "North Star," which is the series "High Desert".
F.2 2WikimQA
@@
MA-RAG 70B Plan: ["Identify the publisher of ’Woman’s Century’.", "Determine the founding date of the identified publisher."]
@@
Extracted Information: doc-12413249: [The National Council of Women of Canada (NCWC) was founded on October 27, 1893, in Toronto, Ontario. This is relevant as it directly answers the question regarding the founding date of the organization.], doc-12413257: [No related information from this document.], doc-17476996: [The National Council of Women of Canada (NCWC) was established in 1894, as indicated by the passage stating, "the Local Council was established in 1894 as the local chapter of the National Council of Women of Canada (NCWC)."], doc-6878896: [The National Council of Women of Canada was founded in 1893. This information is directly stated in the passage.], doc-20577251: [No related information from this document]
@@
Llama3-70B Answer: The company that published Woman’s Century, initially called the ’Woman’s Century Club’, later Woman’s Century Magazine or just ’Woman’s Century’, is uncertain but it’s known that Woman’s Century Club itself predates the magazine as the club was formed in 1891, in Jacksonville, Florida."
@@
Doc-18191576: Citations Sources Woman’s Century Woman’s Century was the official organ of the National Council of Women of Canada (NCWC), published in Toronto between 1913 and 1921. The magazine was relatively conservative, and supported World War I, but stood up for women’s rights and universal suffrage. The founder of "Woman’s Century" was Jessie Campbell MacIver. She had come to Canada from Scotland with her husband, a lawyer, and five children. She became involved in the National Council of Women. The first issue of "Woman’s Century" appeared in May 1913. It was largely produced out of MacIver’s home, with the help of
@@
Doc-18191569: Woman’s Century Woman’s Century was the official organ of the National Council of Women of Canada (NCWC), published in Toronto between 1913 and 1921. The magazine was relatively conservative, and supported World War I, but stood up for women’s rights and universal suffrage. The founder of "Woman’s Century" was Jessie Campbell MacIver. She had come to Canada from Scotland with her husband, a lawyer, and five children. She became involved in the National Council of Women. The first issue of "Woman’s Century" appeared in May 1913. It was largely produced out of MacIver’s home, with the help of her husband
@@
Doc-3964891: The Century Magazine The Century Magazine was first published in the United States in 1881 by The Century Company of New York City, which had been bought in that year by Roswell Smith and renamed by him after the Century Association. It was the successor of "Scribner’s Monthly Magazine" and ceased publication in 1930. In 1921, "The Cambridge History of English and American Literature" summarized the early history of the magazine: The magazine was very successful during the 19th century, most notably for the aforementioned series of articles about the American Civil War, which ran for three years during the
@@
Doc-12413254: also authorized through race-based arguments that linked White women’s enfranchisement to the need to protect the nation from "racial degeneration." Between 1914 and 1921 the NCWC published the monthly journal "Woman’s Century". The purpose was to educate women about public issues and the reforms that were needed, and to provide a forum for discussion by different women’s groups. The title page described it as "A journal of education and progress for Canadian women." The monthly journal was modeled on successful British and American feminist periodicals. It was one of the very few women’s rights journals published in Canada. In 1918,
@@
Doc-18227076: eight-page weekly paper created and managed by the Women’s Centennial Executive Committee of Philadelphia. It was "devoted to the Industrial interest of women" by promoting choice, equal pay for equal work, and greater financial and social autonomy. The New Century Club was a woman’s organization that was established in 1877 to improve the lives of women. It had committees for working women, municipal affairs and self-education. Rather than portraying the opinions of the "radical" viewpoints of some of its members, the Hallowell, its first president, said that they only "whispered… [the] logic of suffrage." As the organization evolved, the moved
@@
AI Messages: Retrieve(query="Woman’s Century publication company founding date")
@@
Doc-18191576: Citations Sources Woman’s Century Woman’s Century was the official organ of the National Council of Women of Canada (NCWC), published in Toronto between 1913 and 1921. The magazine was relatively conservative, and supported World War I, but stood up for womenś rights and universal suffrage. The founder of "Woman’s Century" was Jessie Campbell MacIver. She had come to Canada from Scotland with her husband, a lawyer, and five children. She became involved in the National Council of Women. The first issue of "Woman’s Century" appeared in May 1913. It was largely produced out of MacIver’s home, with the help of
@@
Doc-18191569: Woman’s Century Woman’s Century was the official organ of the National Council of Women of Canada (NCWC), published in Toronto between 1913 and 1921. The magazine was relatively conservative, and supported World War I, but stood up for women’s rights and universal suffrage. The founder of "Woman’s Century" was Jessie Campbell MacIver. She had come to Canada from Scotland with her husband, a lawyer, and five children. She became involved in the National Council of Women. The first issue of "Woman’s Century" appeared in May 1913. It was largely produced out of MacIver’s home, with the help of her husband
@@
Doc-3964891: The Century Magazine The Century Magazine was first published in the United States in 1881 by The Century Company of New York City, which had been bought in that year by Roswell Smith and renamed by him after the Century Association. It was the successor of "Scribner’s Monthly Magazine" and ceased publication in 1930. In 1921, "The Cambridge History of English and American Literature" summarized the early history of the magazine: The magazine was very successful during the 19th century, most notably for the aforementioned series of articles about the American Civil War, which ran for three years during the
@@
Doc-18227076: eight-page weekly paper created and managed by the Women’s Centennial Executive Committee of Philadelphia. It was "devoted to the Industrial interest of women" by promoting choice, equal pay for equal work, and greater financial and social autonomy. The New Century Club was a woman’s organization that was established in 1877 to improve the lives of women. It had committees for working women, municipal affairs and self-education. Rather than portraying the opinions of the "radical" viewpoints of some of its members, the Hallowell, its first president, said that they only "whispered… [the] logic of suffrage." As the organization evolved, the moved
@@
Doc-12741329: in 1847 and the Philadelphia Female Anti-Slavery Society in the 1850s. She helped found the Pennsylvania Woman Suffrage Association in 1869, and she was its first corresponding secretary. At the Philadelphia Centennial Exposition of 1876, Turner was a leader of the Women’s Congress and distributed the newspaper "New Century for Women" that she wrote and edited at the Women’s Pavilion. The New Century Club women’s club was founded in Philadelphia in 1877 following a stirring paper that Turner delivered at the Women’s Congress. Turner was the president from 1879 to 1881 and the first corresponding secretary of the literary, social

### v2

@@ 5 Conclusion
We present MA-RAG, a novel multi-agent framework for Retrieval-Augmented Generation that recasts the RAG pipeline as a coordinated reasoning process among specialized agents. Each agent, responsible for planning, retrieval, extraction, or generation, uses chain-of-thought prompting to enable structured and context-sensitive query resolution. Unlike prior work that tunes individual components or requires supervised training, MA-RAG is entirely training-free and generalizes well across domains and question types. Empirical results on five open-domain and multi-hop QA benchmarks show that MA-RAG outperforms strong LLMs and state-of-the-art RAG baselines, achieving new best results on several datasets. Ablation studies confirm the importance of the planner and extractor: the former decomposes complex queries, while the latter improves retrieval precision. Strategic allocation of model capacity across agents yields further gains in both performance and efficiency. Together, these findings highlight the potential of modular, agent-based reasoning as a scalable and adaptable approach to improving retrieval-augmented generation.
Appendix A Appendix
A.1 Details of Experimental Dataset
A.1.1 Main Experiments
@@ A.2 Experiment Results
A.3 MA-RAG with Internet Access
To further evaluate the capabilities of MA-RAG in practical information-seeking scenarios, we conducted additional experiments by granting MA-RAG access to real-time web search. Specifically, we integrated Google Search as the external retrieval engine and evaluated on the SimpleQA benchmark (Wei et al., 2024). This dataset is designed to assess factual question-answering abilities of frontier models without access to the web. It contains questions that often require up-to-date or obscure knowledge, making it a challenging benchmark for retrieval-augmented systems. We enabled MA-RAG to retrieve evidence via Google Search, replacing the retrieval tool with this search engine.
Results. As shown in Table 6, MA-RAG achieved an accuracy of 86.4% on SimpleQA, demonstrating a significant improvement over GPT-4o (40.1%) and the GPT-4o-mini baseline when equipped with web access. While DeepSeek-R1 (82.4%) performs significantly better than GPT-4o alone, our MA-RAG (GPT-4o-mini, web) results show that multi-agent reasoning substantially boosts performance. Although our current results are still below ODS-v2/ODS-v1+DeepSeek-R1* (Alzubi et al., 2025b), MA-RAG outperforms ODS-v1+Llama3.1-70B (Alzubi et al., 2025b). We believe that the performance gap largely depends on the reasoning model used. Additionally, we highlight that Perplexity Deep Research is a closed-source system without an arXiv paper or public report. These findings suggest that MA-RAG’s combination of multi-agent reasoning and web integration is also highly effective for open-domain factual QA.
Table 6: SimpleQA results (Accuracy) for recent systems. * indicates models allowed to access the internet.
Method SimpleQA (%)
Qwen 2.5 9.1
Llama3.1-70B 20.4
Claude 3.5 Sonnet 28.9
GPT-4o 40.1
DeepSeek-R1 82.4
Perplexity Deep Research 93.9
ODS-v1+Llama3.1-70B 83.4
ODS-v2+DeepSeek-R1 88.3
ODS-v1+DeepSeek-R1 87.7
MA-RAG (GPT-4o-mini, web) 86.4
Note: MA-RAG here is evaluated with live Google Search. All other results are as reported in prior work or reproduced using public checkpoints.
A.4 Implementation Details
(a) MA-RAG graph representations in Langchain.
@@ A.4 Implementation Details
A.5 Workflows
We implement MA-RAG using LangChain and LangGraph [^3] https://www.langchain.com/, where agents communicate through structured JSON messages. Each agent is represented as a node in the graph, and edges determine the next agent to execute based on the current state. The overall graph structure representing the architecture of MA-RAG is shown in Figure 8(a). For modularity and clarity, we define separate subgraphs for core components, including the planner, the plan executor, and the RAG pipeline.
A.5.1 Agent Communication
@@ A.6 Boarder Impacts
A.7 Prompt Formats
A.7.1 Planner Agent
System: You are tasked with assisting users in generating structured plans for answering questions. Your goal is to deconstruct a query into manageable, simpler components. For each question, perform these tasks: *Analysis: Identify the core components of the question, emphasizing the key elements and context needed for a comprehensive understanding. Determine whether the question is straightforward or requires multiple steps to provide an accurate answer. *Plan Creation: - Break down the question into smaller, simpler questions by reasoning that lead to the final answer. Ensure those steps are non overlap. - Ensure each step is clear and logically sequenced. - Each step is a question to search, or to aggregate output from previous steps. Do not verify previous step. Notes: - Put your output in a list of string, each string describe a sub-task User: {Question}
A.7.2 Step Definer Agent
System: Given a plan, the current step, and the results from finished steps, decide the task for this step.Output the type of task and the query. The query need to be in detail, include all of information from previous step’s results in the query if it maked, especially for aggregate task. Be concise. User: Plan: {plan} Current step: {cur_step} Results of finished steps: {memory}
A.7.3 Extractor Agent
System: Summarize and extract all relevant information from the provided passages based on the given question. Remove all irrelevant information. Think step-by-step. **Identify Key Elements**: Read the question carefully to determine what specific information is being requested. **Analyze Passages**: Review the passages thoroughly to find any segments that contain information relevant to the question. **Extract Relevant Information**: Highlight or note down sentences, phrases, or words from the passages that relate to the question. **Remove Irrelevant Details**: Ensure that all extracted information is relevant to the question, eliminating any unnecessary or unrelated content. Output Format - Output a list of notes. Each note contains related information from the passage, and each note is clear, standalone. Notes - Avoiding any irrelevant details. - If a piece of information is mentioned in multiple places, include it only once. - If there are no related information, output: No related information from this document. User: Passage: passage Query: question?
A.7.4 Question-Answering Agent
System: You are an assistant for question-answering tasks. Use the following process to deliver concise and precise answers based on the retrieved context. 1. Analyze Carefully: Begin by thoroughly analyzing both the question and the provided context. 2. Identify Core Details: Focus on identifying the essential names, terms, or details that directly answer the question. Disregard any irrelevant information. 3. Provide a Concise Answer: - Remove redundant words and extraneous details. - Present the answer by listing only the necessary names, terms, or very brief facts that are crucial for answering the question. 4. Clarity and Accuracy: Ensure that your answer is clear and maintains the original meaning of the information provided. 5. Consensus: If the contexts are not consensus, pick one which is the most logical, consensus, or confident. User: Retrieved information: context Question: question
A.8 Case Study
A.8.1 HotpotQA
@@
Q: What role did Thomas Doherty play in sitcom based on the Israeli series ”North Star”? Correct Answer: Sean
MA-RAG 70B Plan: [”Identify the sitcom based on the Israeli series North Star.”, ”Determine the character role played by Thomas Doherty in that sitcom.”]
@@
Extracted Information: doc-19587422: [No related information from this document.], doc-19587424: [No related information from this document.], doc-14539692: [No related information from this document.], doc-19587423: [The sitcom based on the Israeli series ’North Star’ is titled ”The Lodge.” This is evidenced by the passage stating, ”The British Disney Channel produces an adapted under the title ’The Lodge’ (working title: ’North Star’)”.], doc-17716638: [No related information from this document.]
@@
Llama3-70B Answer: Thomas Doherty played the role of Jasper in the sitcom ”High Fidelity”, which is based on the Israeli series ”North Star.”
ChatQA 70B Query: What role did Thomas Doherty play in sitcom based on the Israeli series ”North Star”?
@@
Doc-20322850: Thomas Doherty (actor) Thomas Doherty (born 21 April 1995) is a Scottish actor, best known for his roles as Harry Hook in Disney Channel’s ”Descendants 2” and as Sean in the British musical comedy-drama, ”The Lodge”. Doherty was born and raised in Edinburgh, Scotland. He has an older brother and younger sister, and his parents were both in the banking business. Doherty began acting at the age of 5 and also played football. Doherty joined Lothian Youth Arts and Musicals Company and played a variety of roles, including his debut as Disney’s Aladdin. Doherty attended Royal High School in Edinburgh.
@@
Doc-12942841: them. The show was canceled after one season owing to poor ratings. She also starred in the popular British sitcom ””Bo! in the USA””, a brain child of Leigh Francis. In the show, she plays herself being randomly harassed by Avid Merrion (Francis), who claims they are lovers. The show aired in October 2006 on the British Channel 4, and she appeared in several episodes. During 2007, Doherty appeared in two television projects. She first appeared in ””Kiss Me Deadly: A Jacob Keane Assignment”” for the Here TV network and followed up with a starring role in the holiday film
Doc-1423761: to the stage, making a disastrous start in John Phillips’ ””Man on the Moon”” (1975). In 1977, he returned to his birthplace, Halifax, Nova Scotia, and started playing legitimate roles, including Shakespeare, at the Neptune Theatre under the tutelage of John Neville. This led to television work, beginning with a variety program, ””Denny’s Sho*””, which ran for one season in 1978. He went on to host and voice parts in the children’s program, ””Theodore Tugboat””, and to act in various series, including twenty-two episodes of the drama ””Pit Pony””. Doherty also performed with the New Mamas and the Papas (see
@@
ReAct (70B) Query: What role did Thomas Doherty play in sitcom based on the Israeli series ”North Star”?
AI Messages: Retrieve(query=”Thomas Doherty sitcom North Star role”)
@@
Doc-20322850: Thomas Doherty (actor) Thomas Doherty (born 21 April 1995) is a Scottish actor, best known for his roles as Harry Hook in Disney Channel’s ”Descendants 2” and as Sean in the British musical comedy-drama, ”The Lodge”. Doherty was born and raised in Edinburgh, Scotland. He has an older brother and younger sister, and his parents were both in the banking business. Doherty began acting at the age of 5 and also played football. Doherty joined Lothian Youth Arts and Musicals Company and played a variety of roles, including his debut as Disney’s Aladdin. Doherty attended Royal High School in Edinburgh.
Doc-1423761: to the stage, making a disastrous start in John Phillips”́”Man on the Moon”” (1975). In 1977, he returned to his birthplace, Halifax, Nova Scotia, and started playing legitimate roles, including Shakespeare, at the Neptune Theatre under the tutelage of John Neville. This led to television work, beginning with a variety program, ””Dennyś Sho*””, which ran for one season in 1978. He went on to host and voice parts in the childrenś program, ””Theodore Tugboat””, and to act in various series, including twenty-two episodes of the drama ””Pit Pony””. Doherty also performed with the New Mamas and the Papas (see
Doc-18484135: After high school, Doherty attended The MGA Academy of Performing Arts in Edinburgh, Scotland. He graduated after studying music theatre at MGA Academy in July 2015 and immediately signed a contract with Olivia Bell Management in London. Doherty was trained in acting, singing, and various types of dance including contemporary, hip hop, jazz, tap and ballet. After graduating from The MGA Academy of Performing Arts in the summer of 2015, Doherty began waiting tables in Edinburgh at Tigerlilly restaurant. On his days off, Doherty split his time between auditioning for ””The Lodge”” and working with the Edinburgh Fringe. Doherty was
@@
Doc-2573069: of Harbour Master, as well as the voice-overs of the characters, in ””Theodore Tugboat””, a CBC Television children’s show chronicling the ””lives”” of vessels in a busy harbour loosely based upon Halifax Harbour. In 1999, he also played Charley McGinnis in 22 episodes of the CBC Television series ””Pit Pony””. In 2004, Doherty appeared on Sharon, Lois Bram’s 25th Anniversary Concert special, ””25 Years of Skinnamarink””, that aired on CBC on January 1, 2004. He sang two songs with the trio: ”California Dreamin” and ”Who Put the Bomp?” One of his last appearances was in the Canadian TV series
Answer: Thomas Doherty played the role of Brandon in the sitcom based on the Israeli series ”North Star,” which is the series ”High Desert”.
A.8.2 2WikimQA
@@
MA-RAG 70B Plan: [”Identify the publisher of ’Woman’s Century’.”, ”Determine the founding date of the identified publisher.”]
@@
Extracted Information: doc-12413249: [The National Council of Women of Canada (NCWC) was founded on October 27, 1893, in Toronto, Ontario. This is relevant as it directly answers the question regarding the founding date of the organization.], doc-12413257: [No related information from this document.], doc-17476996: [The National Council of Women of Canada (NCWC) was established in 1894, as indicated by the passage stating, ”the Local Council was established in 1894 as the local chapter of the National Council of Women of Canada (NCWC).”], doc-6878896: [The National Council of Women of Canada was founded in 1893. This information is directly stated in the passage.], doc-20577251: [No related information from this document]
@@
Llama3-70B Answer: The company that published Woman’s Century, initially called the ’Woman’s Century Club’, later Woman’s Century Magazine or just ’Woman’s Century’, is uncertain but it’s known that Woman’s Century Club itself predates the magazine as the club was formed in 1891, in Jacksonville, Florida.”
@@
Doc-18191576: Citations Sources Woman’s Century Woman’s Century was the official organ of the National Council of Women of Canada (NCWC), published in Toronto between 1913 and 1921. The magazine was relatively conservative, and supported World War I, but stood up for women’s rights and universal suffrage. The founder of ”Woman’s Century” was Jessie Campbell MacIver. She had come to Canada from Scotland with her husband, a lawyer, and five children. She became involved in the National Council of Women. The first issue of ”Woman’s Century” appeared in May 1913. It was largely produced out of MacIver’s home, with the help of
@@
Doc-18191569: Woman’s Century Woman’s Century was the official organ of the National Council of Women of Canada (NCWC), published in Toronto between 1913 and 1921. The magazine was relatively conservative, and supported World War I, but stood up for women’s rights and universal suffrage. The founder of ”Woman’s Century” was Jessie Campbell MacIver. She had come to Canada from Scotland with her husband, a lawyer, and five children. She became involved in the National Council of Women. The first issue of ”Woman’s Century” appeared in May 1913. It was largely produced out of MacIver’s home, with the help of her husband
@@
Doc-3964891: The Century Magazine The Century Magazine was first published in the United States in 1881 by The Century Company of New York City, which had been bought in that year by Roswell Smith and renamed by him after the Century Association. It was the successor of ”Scribner’s Monthly Magazine” and ceased publication in 1930. In 1921, ”The Cambridge History of English and American Literature” summarized the early history of the magazine: The magazine was very successful during the 19th century, most notably for the aforementioned series of articles about the American Civil War, which ran for three years during the
@@
Doc-12413254: also authorized through race-based arguments that linked White women’s enfranchisement to the need to protect the nation from ”racial degeneration.” Between 1914 and 1921 the NCWC published the monthly journal ”Woman’s Century”. The purpose was to educate women about public issues and the reforms that were needed, and to provide a forum for discussion by different women’s groups. The title page described it as ”A journal of education and progress for Canadian women.” The monthly journal was modeled on successful British and American feminist periodicals. It was one of the very few women’s rights journals published in Canada. In 1918,
@@
Doc-18227076: eight-page weekly paper created and managed by the Women’s Centennial Executive Committee of Philadelphia. It was ”devoted to the Industrial interest of women” by promoting choice, equal pay for equal work, and greater financial and social autonomy. The New Century Club was a woman’s organization that was established in 1877 to improve the lives of women. It had committees for working women, municipal affairs and self-education. Rather than portraying the opinions of the ”radical” viewpoints of some of its members, the Hallowell, its first president, said that they only ”whispered… [the] logic of suffrage.” As the organization evolved, the moved
@@
AI Messages: Retrieve(query=”Woman’s Century publication company founding date”)
@@
Doc-18191576: Citations Sources Woman’s Century Woman’s Century was the official organ of the National Council of Women of Canada (NCWC), published in Toronto between 1913 and 1921. The magazine was relatively conservative, and supported World War I, but stood up for womenś rights and universal suffrage. The founder of ”Woman’s Century” was Jessie Campbell MacIver. She had come to Canada from Scotland with her husband, a lawyer, and five children. She became involved in the National Council of Women. The first issue of ”Woman’s Century” appeared in May 1913. It was largely produced out of MacIver’s home, with the help of
@@
Doc-18191569: Woman’s Century Woman’s Century was the official organ of the National Council of Women of Canada (NCWC), published in Toronto between 1913 and 1921. The magazine was relatively conservative, and supported World War I, but stood up for women’s rights and universal suffrage. The founder of ”Woman’s Century” was Jessie Campbell MacIver. She had come to Canada from Scotland with her husband, a lawyer, and five children. She became involved in the National Council of Women. The first issue of ”Woman’s Century” appeared in May 1913. It was largely produced out of MacIver’s home, with the help of her husband
@@
Doc-3964891: The Century Magazine The Century Magazine was first published in the United States in 1881 by The Century Company of New York City, which had been bought in that year by Roswell Smith and renamed by him after the Century Association. It was the successor of ”Scribner’s Monthly Magazine” and ceased publication in 1930. In 1921, ”The Cambridge History of English and American Literature” summarized the early history of the magazine: The magazine was very successful during the 19th century, most notably for the aforementioned series of articles about the American Civil War, which ran for three years during the
@@
Doc-18227076: eight-page weekly paper created and managed by the Women’s Centennial Executive Committee of Philadelphia. It was ”devoted to the Industrial interest of women” by promoting choice, equal pay for equal work, and greater financial and social autonomy. The New Century Club was a woman’s organization that was established in 1877 to improve the lives of women. It had committees for working women, municipal affairs and self-education. Rather than portraying the opinions of the ”radical” viewpoints of some of its members, the Hallowell, its first president, said that they only ”whispered… [the] logic of suffrage.” As the organization evolved, the moved
@@
Doc-12741329: in 1847 and the Philadelphia Female Anti-Slavery Society in the 1850s. She helped found the Pennsylvania Woman Suffrage Association in 1869, and she was its first corresponding secretary. At the Philadelphia Centennial Exposition of 1876, Turner was a leader of the Women’s Congress and distributed the newspaper ”New Century for Women” that she wrote and edited at the Women’s Pavilion. The New Century Club women’s club was founded in Philadelphia in 1877 following a stirring paper that Turner delivered at the Women’s Congress. Turner was the president from 1879 to 1881 and the first corresponding secretary of the literary, social
@@
A.9 LLM Usage Declaration
We declare that large language models (LLMs) were used in limited and specific capacities in this work. Specifically, LLMs such as LLaMA3 (8B and 70B) and GPT-4o-mini were used as components within the MA-RAG framework for retrieval-augmented reasoning. In addition to framework development, LLMs were employed for grammar checking and language refinement of the manuscript. All core technical contributions, experimental design, analysis, and conclusions presented in this work are entirely our own. The use of LLMs did not influence the scientific methodology, results interpretation, or technical contributions of this research.